# Probing Hernquist dark matter with black hole shadows: A comprehensive study of various accretions

### v2

@@ Abstract
The shadow of a black hole is critically dependent on the surrounding accreting matter. We investigate the observational signatures of a Schwarzschild black hole embedded in a Hernquist dark matter (DM) halo under three distinct accretion scenarios: a geometrically thin disk, a static spherical flow, and an infalling spherical flow. For the thin disk model, we find that direct emission dominates the total observed intensity, while the size and brightness of the lensing and photon rings serve as sensitive probes of the Hernquist DM parameters. Under spherical accretion, the Hernquist DM halo significantly enlarges the photon sphere. This results in an observable shadow that is approximately $2\%$ to $30\%$ larger than in the vacuum case, though this increase in size is accompanied by a considerable decrease in overall image brightness. Furthermore, the Doppler de-boosting effect in the infalling scenario produces a markedly darker image than its static counterpart. Our results demonstrate that the size and brightness profile of a black hole shadow provide a powerful observational tool to probe and constrain the distribution of dark matter in galactic centers.
@@ I Introduction
The observation of black hole shadows has entered a new era, establishing them as a ground-breaking tool for researching powerful gravitational forces. The fundamental nature of gravity has been further illuminated by analysis of the critical behaviour of photons close to black holes. Recent advances in the Event Horizon Telescope’s (EHT) imaging of the supermassive black holes $\text{M87}^{*}$ Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023) and $\text{SgrA}^{*}$ Ghez et al. (2005); Boehle et al. (2016); Akiyama et al. (2022a, b, c, d, e, f, 2024a, 2024b) have shown shadow features that not only support general relativity (GR) Hawking (1972); Bardeen et al. (1973a); Wald (1994, 2010) predictions but also provide a novel window of probing black holes. The black hole shadow, also known as the photon capture zone, is the centre dark area shown in these pictures Perlick and Tsupko (2022). This is encircles a compact, asymmetrical and dazzling circular structure. This phenomenon, known as the photon ring, results from light beams from infinity being deflected by the strong gravitational field as they approach the black hole. In backward ray-tracing methods, light rays that approach the critical impact parameter asymptotically converge to the bound photon orbit. The so-called photon sphere, which projects the shadow of the black hole, is made up of this confined photon orbit. An angular radius equation for the shadow of Schwarzschild black holes has been established by Synge and Luminet using a crucial impact parameter of $b_{p}=3\sqrt{3}M$ Synge (1966); Luminet (1979). The first image of a black hole surrounded by a thin accretion disk was analytically produced by the authors in Ref.Luminet (1979), displaying primary and secondary images that emerge beyond the black hole’s shadow. Bardeen later studied the D-shaped shadow of Kerr black holes, showing that the shape of the shadow can be changed by the spin of the black hole Bardeen et al. (1973b); Chandrasekhar (1998). Additionally, a great deal of research has been done on black hole shadows in higher-dimensional spacetime and other modified gravity theories Amarilla et al. (2010); Amarilla and Eiroa (2012); Amir et al. (2018); Mizuno et al. (2018); Eiroa and Sendra (2018); Banerjee et al. (2020); Bambi et al. (2019); Vagnozzi et al. (2023); Khodadi et al. (2024). In contrast to single shadows, the research has also suggested multiple shadows for wormholes and black holes Huang et al. (2016); Guerrero et al. (2021). While the finding of the EHT mainly relates to black holes in GR, it provides a great deal of opportunity to investigate other compact objects in theories other than GR.
@@
To investigate the main characteristics of black hole pictures and find possible indications of novel physics, it is helpful to set up a basic accretion model. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accret ion settings the shadow has resilient properties where its shape and size are determined by spacetime instead of accretion details Gralla et al. (2019); Tsukamoto et al. (2014); Tsukamoto (2018). Further investigations of optically thin and geometrically thin accretion models show that the size of the shadow is strongly related to the position of the accretion Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are always present in the bright areas outside the shadow. These characteristics make it possible to distinguish GR black holes from other compact objects or black holes in modified gravity theories through observation Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are extremely important for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
@@
The environment surrounding the black hole, in particular the distribution of DM, offers an additional convincing explanation for possible departures from conventional predictions, in addition to modified gravity theories. The discovery of huge elliptical and spiral galaxies was the first significant development in the seek for DM Rubin et al. (1980). Persic found that about $90\%$ of a galaxy’s mass is made up of DM Persic et al. (1996). Estimating DM’s contribution close to the galactic centre is especially important since there is strong evidence that DM halos around astrophysical black holes Collaboration (2019a, f); Sofue (2013); Boshkayev and Malafarina (2019); Jusufi et al. (2019); Konoplya (2019); Konoplya and Zhidenko (2022); Hou et al. (2018a); Liang et al. (2023); Gohain et al. (2024); Övgün and Pantig (2025); Jha (2025). Several DM distributions are covered by the Dehnen density distribution, which is frequently used to describe dwarf galaxies Dehnen (1993); Mo et al. (2010). These investigations usually use the $\text{Dehnen}-(1,4,1)$ type, which is the Hernquist DM distribution. The Hernquist decay is more consistent with the dynamical evidence of dwarf elliptical galaxies under observational restrictions. Nevertheless, there is still a dearth of comprehensive investigation into how Hernquist DM halos alter the Schwarzschild black hole shadow’s observable properties. Although the effects of Hernquist DM halos on the photon sphere radius and shadow radius of Schwarzschild black holes are discussed in Ref.Jha (2025); Ahmed et al. (2025); Feng and Zhang (2025), accretion models are not taken into account. Furthermore, the direct contribution of important factors like centre density and core radius to lensing rings and photon rings is not quantified in accretion model studies, which mostly concentrate on vacuum or uniform DM backgrounds Gralla et al. (2019); Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021a); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025); Shaikh et al. (2019); Shi and Cheng (2024); Feng and Zhang (2025); Amir et al. (2018); Eiroa and Sendra (2018); Amarilla and Eiroa (2012); Konoplya (2019); Mizuno et al. (2018); Meng et al. (2023). There is currently no systematic investigation that combines comprehensive accretion models with a Hernquist DM halo. We cannot use black hole shadows as accurate probes to restrict DM characteristics due to this gap.
@@
In this work, we present the first comprehensive characterization of the shadow of Schwarzschild-Hernquist black holes with different accretion models. In order to clarify the associated impacts of Hernquist DM halo parameters on the photon sphere and observed brightness, we thoroughly investigated three thin disk accretion models and two spherical accretion models. Apparently, different accretion regimes produce different shadow pictures under the same conditions. In particular, pictures made up of the shadow, lensing ring, and photon ring are created via thin disk accretion. The shadow and photon ring are spherically symmetric for spherical accretion. The observed intensity is affected differently by static and falling accretion, which significantly changes the brightness and radius of the shadow. A darker shadow than under static conditions results from the extra Doppler effect introduced by falling accretion Li and He (2021b, a); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). We close a gap in the references on Hernquist metrics and accretion models by revealing scaling rules for shadow size and brightness as factors of the Hernquist DM halo distribution. This offers an observable diagnostic standard for differentiating amongst Hernquist DM characteristics.
@@
The structure of this paper is as follows: Sec.II derives the zero-geodesic equation under the Hernquist metric, defining the photon sphere radius and collision parameter; Sec.III analyses the thin disk accretion model, including the contribution regions of direct radiation, lensing rings, and photon rings, and relates the emission intensity to the observed appearance; Sec.IV examines static and falling spherical accretion models, contrasting how varying Hernquist DM halo parameters affect shadow brightness; finally, Sec.V synthesises the findings, emphasising the potential interpretative power of Hernquist DM halos for shadows and outlining their prospects for detecting DM distributions in galactic centres.
@@ II Light deflection and photon orbit
where $\rho_{c}$ and $r_{s}$ represent the center density and core radius of the DM halo, respectively. At short radii $(r\ll r_{s})$, this distribution shows an isothermal feature of $\rho\propto r^{-1}$, but at high radii $(r\gg r_{s})$, it decays to $\rho\propto r^{-4}$Jha (2025). For details of the computational process, we refer the reader to Refs.Mo et al. (2010); Yang et al. (2024); Xu et al. (2018). Here, the metric of a Schwarzschild black hole immersed in the Hernquist DM halo surroundings can be written as follows Jha (2025); Ahmed et al. (2025); Feng and Zhang (2025):
@@
where
@@
Here, the Newton constant $G$ has been set to $1$ for simplicity. Naturally, this black hole metric (3) reduces to the Schwarzschild black hole in the limit $\rho_{c}\to 0$ or $r_{s}\to 0$. Our goal is to investigate the Schwarzschild-Hernquist black hole’s null geodesic. We restrict our analysis to the equatorial plane, $\theta=\pi/2$ and $\dot{\theta}=0$Synge (1966); Bardeen et al. (1972); Gralla and Lupsasca (2020). Therefore, the Lagrangian $\mathcal{L}$ for a particle in this spacetime can be easily obtained,
$\mathcal{L}$ $=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}$ (4)
@@
where, $\dot{x}^{\mu}=\partial x^{\mu}/\partial\chi$ is the four-velocity of the photon and $\chi$ is the affine parameter. Since the coefficient of the metric equation cannot be directly found using the $t$ and $\theta$ coordinates, energy and angular momentum are represented by the conserved variables $E$ and $L$, respectively. That’s Shaikh et al. (2019); He et al. (2022); Li and He (2021a); Shi and Cheng (2024); Feng and Zhang (2025)
@@
By taking into account that $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$ for the null geodesic, and then using Eqs.(6)-(7) to solve Eq.(4), it produces Amir et al. (2018); Eiroa and Sendra (2018); Shaikh et al. (2019); He et al. (2022); Li and He (2021a)
@@
This equation shows that when the photon sphere radius $r_{p}$ changes, the black hole’s effective potential reaches a maximum, which corresponds to real and positive roots. Consequently, the connection between the critical impact parameter $b_{p}$ and the photon sphere radius $r_{p}$ is as followsGralla and Lupsasca (2020); Li and He (2021b); Fathi and Cruz (2023); He et al. (2022); Shi and Cheng (2024):
@@
Figure 1: The effective potential $V_{\text{ph}}$ and the impact parameter $b_{p}$ of the Schwarzschild-Hernquist black holes.
@@
Figure 2: The behaviour of light trajectories under various $\rho_{c}$ and $r_{s}$ in the polar coordinate system $(r,\varphi)$.

### v3

@@ Abstract
The shadow of a black hole is critically dependent on the surrounding accreting matter. We perform a systematic comparative analysis of the observational signatures of a Schwarzschild black hole embedded in a Hernquist dark matter (DM) halo under three distinct accretion scenarios: a geometrically thin disk, a static spherical flow, and an infalling spherical flow. For the thin disk model, we find that direct emission dominates the total observed intensity, while the size and brightness of the lensing and photon rings serve as sensitive probes of the Hernquist DM parameters. From a geometric perspective, the Hernquist DM halo significantly enlarges the photon sphere, resulting in an observable shadow radius approximately $2\%$ to $30\%$ larger than in the vacuum case. Regarding the radiative signatures, the measured intensity profiles, which rely on the particular accretion models, show a general brightness suppression, which is especially affected by the Doppler de-boosting in the infalling scenario. Our results suggest that the size and brightness profile of a black hole shadow provide a valuable theoretical framework for constraining the distribution of dark matter in galactic centers.
@@ I Introduction
The observation of black hole shadows has entered a new era, establishing them as a ground-breaking tool for probing strong gravity regimes. The fundamental nature of gravity has been further illuminated by analysing the critical behaviour of photons close to black holes. Recent advances in the Event Horizon Telescope’s (EHT) imaging of the supermassive black holes $\text{M87}^{*}$ Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023) and $\text{SgrA}^{*}$ Ghez et al. (2005); Boehle et al. (2016); Akiyama et al. (2022a, b, c, d, e, f, 2024a, 2024b) have shown shadow features that not only support general relativity (GR) Hawking (1972); Bardeen et al. (1973); Wald (1994, 2010) predictions but also provide a novel window for probing black holes. The black hole shadow, also known as the photon capture zone, is the central dark area observed in these images Perlick and Tsupko (2022). It is encircled by a compact, asymmetrical and bright circular structure. This phenomenon, known as the photon ring, results from light rays from infinity being deflected by the strong gravitational field as they approach the black hole. In backward ray-tracing methods, light rays that approach the critical impact parameter asymptotically converge to the bound photon orbit. The so-called photon sphere, which projects the shadow of the black hole, consists of this confined photon orbit. An angular radius equation for the shadow of Schwarzschild black holes has been established by Synge and Luminet using a critical impact parameter of $b_{p}=3\sqrt{3}M$ Synge (1966); Luminet (1979). The first image of a black hole surrounded by a thin accretion disk was analytically produced by the authors in Ref.Luminet (1979), displaying primary and secondary images that appear outside the black hole’s shadow. Bardeen later studied the D-shaped shadow of Kerr black holes, showing that the shape of the shadow can be changed by the spin of the black hole Bardeen and others (1973); Chandrasekhar (1998). Additionally, extensive research has been conducted on black hole shadows in higher-dimensional spacetime and other modified gravity theories Amarilla et al. (2010); Amarilla and Eiroa (2012); Amir et al. (2018); Mizuno et al. (2018); Eiroa and Sendra (2018); Banerjee et al. (2020); Bambi et al. (2019); Vagnozzi et al. (2023); Khodadi et al. (2024). In contrast to single shadows, studies have also suggested multiple shadows for wormholes and black holes Huang et al. (2016); Guerrero et al. (2021). While the findings of the EHT mainly relate to black holes in GR, they provide significant opportunities to investigate other compact objects in theories other than GR.
@@
To investigate the observational characteristics of black hole images and search for potential indications of novel physics, it is essential to employ simplified accretion models. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accretion settings the shadow has robust properties, where its shape and size are determined by spacetime geometry rather than accretion details Gralla et al. (2019); Tsukamoto et al. (2014); Tsukamoto (2018). Conversely, investigations of optically thin and geometrically thin accretion models show that the appearance of the shadow is strongly related to the position of the emission source Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are the three main components that these models usually appear as in the seen image. These distinct signatures offer a potential means to distinguish GR black holes from other compact objects or black holes in modified gravity theories via observational data Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are crucial for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
@@
The environment surrounding the black hole, in particular the distribution of DM, offers a compelling explanation for possible departures from conventional predictions, in addition to modified gravity theories. The discovery of massive elliptical and spiral galaxies marked a significant development in the search for DM Rubin et al. (1980). Succeeding studies indicated that about $90\%$ of a galaxy’s mass is composed of DM Persic et al. (1996). Estimating DM’s contribution close to the galactic centre is especially important since there is strong evidence that DM halos around astrophysical black holes Collaboration (2019a, f); Sofue (2013); Boshkayev and Malafarina (2019); Jusufi et al. (2019); Konoplya (2019); Konoplya and Zhidenko (2022); Hou et al. (2018b); Liang et al. (2023); Gohain et al. (2024); Övgün and Pantig (2025); Jha (2025). Several DM distributions are covered by the Dehnen density distribution, which is frequently used to describe dwarf galaxies Dehnen (1993); Mo et al. (2010). These investigations usually use the $\text{Dehnen}-(1,4,1)$ type, corresponding to the Hernquist DM distribution, is widely used to describe elliptical galaxies. The Hernquist profile provides a mathematically tractable description that is consistent with the dynamical evidence of galaxies. Nevertheless, more detailed examination is necessary to understand the distinct observational signatures of Schwarzschild black holes embedded in Hernquist DM halos, especially under realistic accretion scenarios. Although the effects of Hernquist DM halos on the photon sphere radius and shadow radius have been discussed in Ref.Jha (2025); Ahmed et al. (2025); Feng and Zhang (2025), accretion models remains less explored. Furthermore, quantitative studies linking halo parameters, such as central density and core radius, to the structure of lensing and photon rings are limited, as most existing works concentrate on vacuum or uniform DM backgrounds Gralla et al. (2019); Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025); Shaikh et al. (2019); Shi and Cheng (2024); Feng and Zhang (2025); Amir et al. (2018); Eiroa and Sendra (2018); Amarilla and Eiroa (2012); Konoplya (2019); Mizuno et al. (2018); Meng et al. (2023). This work aims to bridge this gap by conducting a systematic comparison of accretion models within a Hernquist DM halo. Addressing this gap is essential for using black hole shadows as accurate probes to restrict DM characteristics.
@@
In this paper, we present a systematic comparative study of the shadow of Schwarzschild-Hernquist black holes with different accretion models. In order to clarify the associated impacts of Hernquist DM halo parameters on the photon sphere and observed brightness, we thoroughly investigated three thin disk accretion models and two spherical accretion models. Our results demonstrate that different accretion regimes produce different shadow images under the same background geometry. Specifically, pictures made up of the shadow, lensing ring, and photon ring are created via thin disk accretion. The shadow and photon ring are spherically symmetric for spherical accretion. A darker shadow than under static conditions results from the extra Doppler effect introduced by falling accretion Li and He (2021a, b); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). We aim to close a gap in the references on Hernquist metrics and accretion models by revealing scaling rules for shadow size and brightness as functions of the Hernquist DM halo distribution. This offers a potential diagnostic framework for constraining Hernquist DM characteristics.
@@
The structure of this paper is as follows: Sec.II derives the null geodesic under the Hernquist metric, defining the photon sphere radius and critical impact; Sec.III analyses the thin disk accretion model, including the contribution regions of direct radiation, lensing rings, and photon rings, and relates the emission intensity to the observed appearance; Sec.IV examines static and infalling spherical accretion models, contrasting how varying Hernquist DM halo parameters affect shadow brightness; finally, Sec.V summarizes the findings, emphasising the potential interpretative power of Hernquist DM halos for shadows and outlining their prospects for detecting DM distributions in galactic centres.
@@ II Light deflection and photon orbit
where $\rho_{c}$ and $r_{s}$ represent the center density and core radius of the DM halo, respectively. The Hernquist DM mass profile can be expressed as follows,
$M_{H}=\int_{0}^{r}4\pi\rho(r^{\prime})r^{\prime 2}\mathrm{d}r^{\prime}=\dfrac{2\pi r^{2}r_{s}^{3}\rho_{s}}{(r+r_{s})^{2}}.$ (2)
According to the derivation in Ref.Jha (2025), the spherically symmetric metric for the combined system of black hole and Hernquist DM halo may be assumed,
@@
where the metric function is the linear superposition of the two,
$f(r)=f_{\text{H}}(r)+f_{\text{BH}}(r).$ (4)
The tangential velocity of test particles $v_{t}$ and the mass profile $M_{\text{H}}$ in the halo determine the redshift function $f_{\text{H}}(r)$ via the relation Jha (2025); Ahmed et al. (2025),
$v_{t}^{2}=\dfrac{M_{\text{H}}}{r}=r\dfrac{\mathrm{d}}{\mathrm{d}r}\ln\sqrt{f_{\text{H}}(r)}.$ (5)
The redshift function $f_{\text{H}}(r)$ takes on an exponential shape when this relation is integrated. Assuming the halo potential is modest in relation (5) to the central black hole term, a linear approximation is made to the halo contribution to generate the analytically tractable form,
$f_{\text{H}}(r)$ $=\exp\left(-\dfrac{4\pi\rho_{s}r_{s}^{3}}{r+r_{s}}\right)$
$\simeq 1-\dfrac{4\pi\rho_{s}r_{s}^{3}}{r+r_{s}}.$ (6)
Importantly, even though this metric was obtained using the aforementioned approximation, it is an accurate solution to the Einstein field equations in terms of its physical status Jha (2025),
$R_{\mu\nu}-\dfrac{1}{2}Rg_{\mu\nu}=\kappa^{2}\left(T_{\mu\nu}^{\text{H}}+T_{\mu\nu}^{\text{B}}\right),$ (7)
The energy-momentum tensor corresponds to an anisotropic fluid for the redshift function (II) in the line element (10) might be expressed as
$T_{\mu}^{\nu}=g^{\nu\delta}T_{\mu\delta}=\text{diag}\left[-\rho,p_{r},p,p\right].$ (8)
The Schwarzschild black hole’s energy-momentum tensor $T_{\mu\nu}^{B}=0$ since it is a vacuum solution. As a result, the dark matter tensor $T_{\mu\nu}^{\text{H}}$ dominates the source term of the complex system. This means that $f_{\text{BH}}(r)$ must be reduced to the Schwarzschild solution in the vacuum in order for the equations to be solved Jha (2025); Ahmed et al. (2025),
$f_{\text{BH}}(r)=-\dfrac{2M}{r}.$ (9)
At short radii $(r\ll r_{s})$, this distribution shows an isothermal feature of $\rho\propto r^{-1}$, but at high radii $(r\gg r_{s})$, it decays to $\rho\propto r^{-4}$Jha (2025). For details of the computational process, we refer the reader to Refs.Mo et al. (2010); Yang et al. (2024); Xu et al. (2018). Here, the metric of a Schwarzschild black hole immersed in the Hernquist DM halo surroundings can be written as follows Jha (2025); Ahmed et al. (2025); Feng and Zhang (2025):
@@
Figure 1: The metric function $f(r)$ as a function of the radial coordinate $r/M$ for different Hernquist DM halo parameters. The horizontal line represents $f(r)=0$. A close-up of the area close to the event horizon is shown in the inset. In contrast to the Schwarzschild case (black line), it is clear that the zero-crossing point, which determines the event horizon radius $r_{h}$, spreads outward as the halo parameters $\rho_{c}$ and $r_{s}$ grow.
Here, the Newton constant $G$ has been set to $1$ for simplicity. Naturally, this black hole metric (10) reduces to the Schwarzschild black hole in the limit $\rho_{c}\to 0$ or $r_{s}\to 0$. In Fig.1, we depict the behavior of the metric function $f(r)$ to show how the Hernquist DM halo affects spacetime. The largest root of the equation $f(r_{h})=0$ defines the event horizon radius $r_{h}$. This critical point in the Schwarzschild scenario is located at $r_{h}=2M$. However, the DM halo causes the curve to cross the zero axis at a wider radius, as seen in Fig.1. In particular, a monotonic expansion of the event horizon is induced by an increase in either the core radius $r_{s}$ or the central density $\rho_{c}$. This geometric change indicates that the input from the DM halo may successfully boost the black hole’s.
Our goal is to investigate the Schwarzschild-Hernquist black hole’s null geodesic. We restrict our analysis to the equatorial plane, $\theta=\pi/2$ and $\dot{\theta}=0$ Synge (1966); Bardeen et al. (1972); Gralla and Lupsasca (2020). Therefore, the Lagrangian $\mathcal{L}$ for a particle in this spacetime can be easily obtained,
$\mathcal{L}$ $=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}$
@@
where, $\dot{x}^{\mu}=\partial x^{\mu}/\partial\chi$ is the four-velocity of the photon and $\chi$ is the affine parameter. Since the coefficient of the metric equation cannot be directly found using the $t$ and $\theta$ coordinates, energy and angular momentum are represented by the conserved variables $E$ and $L$, respectively. That’s Shaikh et al. (2019); He et al. (2022); Li and He (2021b); Shi and Cheng (2024); Feng and Zhang (2025)
@@
By taking into account that $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$ for the null geodesic, and then using Eqs.(12)-(13) to solve Eq.(II), it produces Amir et al. (2018); Eiroa and Sendra (2018); Shaikh et al. (2019); He et al. (2022); Li and He (2021b)
@@
This equation shows that when the photon sphere radius $r_{p}$ changes, the black hole’s effective potential reaches a maximum, which corresponds to real and positive roots. Consequently, the connection between the critical impact parameter $b_{p}$ and the photon sphere radius $r_{p}$ is as followsGralla and Lupsasca (2020); Li and He (2021a); Fathi and Cruz (2023); He et al. (2022); Shi and Cheng (2024):
@@
Figure 2: The effective potential $V_{\text{ph}}$ as a function of the radius $r/M$ for the Schwarzschild-Hernquist black holes with varying Hernquist DM parameters. We set $M=1$ here. The peak of each curve corresponds to the photon sphere radius $r_{ph}$, which determines the critical impact parameter $b_{p}$. Three different physical regions of the photon trajectories are distinguished by the critical value: scattering to infinity (Region 1, $b>b_{p}$), critical unstable orbit (Region 2, $b=b_{p}$) and capture (Region 3, $b<b_{p}$).
@@
Figure 3: The behaviour of light trajectories under various $\rho_{c}$ and $r_{s}$ in the polar coordinate system $(r,\varphi)$. The solid black disk and the green dashed circle indicates the event horizon and the photon shpere, respectively. The yellow lines correspond to scattered rays $(b>b_{p})$, the blue lines to captured light $(b<b_{p})$ and the red one means the critical unstable orbit $(b=b_{p})$.